An electric field E can pull an electron out of a Helium atom in the same time that the electron takes
to go around the nucleus once. Then E is?
1
Expert's answer
2020-05-28T11:48:58-0400

It will be possible when force produced by the electric field will be equal to the force produced by the electrostatic force of attraction between nucleus and electron

"eE=\frac{kZe^2}{r^2}\\E=\frac{k.4e}{r^2}"